wein's displacement law fails at


a)high temperature


b)low temperature


c)short wavelength


d)long wavelength

Grade:

3 Answers

Aman Bansal
12 years ago

Dear Harsh,

Wien s displacement law states that the wavelength distribution of thermal radiation from a black body at any temperature has essentially the same shape as the distribution at any other temperature, except that each wavelength is displaced on the graph. Apart from an overall T3multiplicative factor, the average thermal energy in each mode with frequency ν only depends on the ratio ν / T. Restated in terms of the wavelength λ = c / ν, the distributions at corresponding wavelengths are related, where corresponding wavelengths are at locations proportional to 1 / T. Blackbody radiation approximates to Wien s law at high frequencies.

From this general law, it follows that there is an inverse relationship between the wavelength of the peak of the emission of a black body and itstemperature when expressed as a function of wavelength, and this less powerful consequence is often also called Wien s displacement law in many textbooks.

\lambda_\max = \frac{b}{T}

where λmax is the peak wavelength, T is the absolute temperature of the black body, and b is a constant of proportionality called Wien s displacement constant, equal to 2.8977685(51)×10−3 m·K (2002 CODATA recommended value).

For wavelengths near the visible spectrum, it is often more convenient to use the nanometer in place of the meter as the unit of measure. In this case, b = 2,897,768.5(51) nm·K.

In the field of plasma physics temperatures are often measured in units of electron volts and the proportionality constant becomesb = 249.71066 nm·eV.

Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Jit Mitra
12 years ago

Short Wavelength ---- Weins law applicable

Long Wavelength  ---- Raleigh Jeans Law

rhea
6 years ago
In cyclotron time period of revolution inside Dees a)increases with radius of circle b)decrease with radius of circle c)is constant d)unpredictable